the perimeter of a retangle photgraph is 35cm. The length is 4cm greater then the width. solve for the length and the width.
Answer by stanbon(75887)

the perimeter of a retangle photgraph is 35cm. The length is 4cm greater then the width. solve for the length and the width.
--------------------
P = 2(length + width)
35 = 2((w+4)+w)
35 = 2(2w+8)
35 = 4w + 16
4w = 19
w = 4 3/4 cm (width)
length = 4 3/4 + 4 = 8 3/4 cm
===================================
Cheers,
Stan H.
===============